# Conefield approach to identifying regions without flux surfaces for magnetic fields

### v1

@@ Abstract
The conefield variant of a Converse KAM method for 3D vector fields, identifying regions through which no invariant 2-tori pass transverse to a specified direction field, is tested on some helical perturbations of an axisymmetric magnetic field in toroidal geometry. This implementation computes bounds on the slopes of invariant tori of a given class and allows to apply a subsidiary criterion for the extension of the non-existence region, saving significant computation time. The method finds regions corresponding to magnetic islands and chaos for the fieldline flow.
Keywords: Magnetic field, Flux surface, Converse KAM method, Conefield
@@ 1 Introduction
The present work presents a new application of Converse KAM theory to magnetic fields, complementing the results of [5]. First, it applies an alternative, yet equivalent, implementation of the method described in [6], called the “conefield method”. This version of the Converse KAM method constructs bounds for the slope of a torus passing through a point when restricted to a transverse section (e.g., poloidal plane). Secondly, as discussed in the following sections, this approach enables processing the method’s output data to further extend the region of non-existence by a subsidiary criterion called “killends”, inspired by [3].
The focus of this study is on the same magnetic fields in toroidal configurations considered in [5], with the same purpose of identifying regions where no invariant tori (flux surfaces) of a given class are present. A class of tori is defined by prescribing a direction field almost everywhere and requiring that the tori be transverse to this field. The primary choice for the direction field is the gradient of an appropriate distance function from a closed field line (magnetic axis), measured with respect to a given metric.
In Section 2, we introduce the magnetic fields analyzed in this paper. Section 3 outlines the application of the Converse KAM method to these magnetic fields and presents the subsidiary criterion to enlarge the detected non-existence region. In Section 4, we present the results of the numerical implementation of the Converse KAM method and the subsidiary criterion. Section 5 presents our conclusions.
@@ 2 Toroidal helical magnetic fields
Given a coordinate system $(x^{1},x^{2},x^{3})$, we will write a magnetic field $B$ in terms of its contravariant components $B^{i}$, rather than its physical ones. This choice reduces the equations of motion for fieldline flow to simply $\dot{x}^{i}=B^{i}(x)$; here “time” is to be understood along the magnetic field lines.
The fields considered have a circular magnetic axis of radius $R_{0}>0$ in the horizontal plane $z=0$. They have a simpler form when described and treated in an adapted toroidal coordinate system $(\psi,\vartheta,\phi)$, which is a variant of the standard toroidal coordinates $(r,\theta,\phi)$. The latter coordinates are related to Cartesian coordinates $(x,y,z)$ through
@@
Following [7], we introduce adapted coordinates $(\psi,\vartheta)$ to make the restriction $\beta_{T}$ of the magnetic flux-form $\beta$ to a poloidal section (ϕ=italic-$\phi=$ constant) take the form
@@
where $\wedge$ denotes the exterior product of differential forms (see [8] for a tutorial). To achieve this, we first define $\psi$ as the toroidal magnetic flux across the poloidal disk of radius $r$ about a point of the magnetic axis, divided by $2\pi$. Thus, integrating
$\beta_{\text{T}}=rRB^{\phi}\,dr\wedge d\theta,$ (2)
over a poloidal disk of radius $r$ yields $\psi$. Then $\vartheta$ can be constructed by equating (1) and (2), up to choice of origin that we take to be at $\theta=0$. For simplicity, we choose our magnetic fields to all have $B^{\phi}={B_{0}R_{0}}/{R^{2}}$, which corresponds to a vertical current along the $z$-axis. Given the previous choices, we arrive at [7, 9, 5]
@@ 2.1 Magnetic fields studied
where $w_{1}\in\mathbb{R}$, $w_{2}\neq 0$, $m,n$ are integers with $m\geq 2$, $f_{mn}$ are smooth functions and $\zeta_{mn}$ arbitrary phases. The factor $\psi^{m/2}$ is to make the resulting vector potential smooth at $\psi=0$.
@@ 3 Converse KAM for magnetic fields
Let $B$ be a nowhere-zero $C^{1}$ vector field on an orientable $3D$ manifold $M$. A torus in $M$ means a homeomorphic image of the standard 2-torus $\mathbb{T}^{2}$, with $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. We define a class of tori in $M$ by specifying a $C^{1}$ vector field $\xi$, called “direction field”, and requiring the tori to be transverse to $\xi$. For differentiable tori, this means that $\xi$ is nowhere tangent to the torus. For non-smooth tori, we require only topological transversality: that the direction field passes from one side of the torus to the other.
In [5], the Converse KAM criterion consisted in pushing forward the direction field $\xi$ along $B$ (in both directions of time) and testing whether it becomes a negative multiple of $\xi$ modulo $B$ at some time. In the present work, we implement a different yet equivalent approach, which we term the ‘conefield formulation’. In this formulation, we pull back a 1-form (in both directions of time) to produce upper and lower bounds on the “slope” of invariant tori through a given point; thus, for each point tested, we construct a cone of possible slopes, forming a “conefield”. If at some stage the upper bound is below the lower bound then the set of possible slopes is empty, so there is no invariant torus through that point.
@@ 3.2 Conefield construction and nonexistence condition
We describe here the idea and construction of conefields for the class of invariant tori of a 3D vector field $B$ that are transverse to $\xi$, by adapting the construction for area-preserving twist maps in [3]. See also the brief description in [6]. We simplify the presentation by restricting attention to differentiable invariant tori.
For the present paper, a cone at a point in 3D is a closed union of 2D subspaces. [^2] This is not the standard definition of a cone but is similar to other non-standard definitions, such as in hyperbolicity theory. A conefield is a field of cones in 3D. A conefield is compatible with a differentiable embedded torus if the tangent plane to the torus at each of its points lies in the conefield.
If there is a differentiable invariant 2-torus $\mathcal{T}$ for $B$ then it is possible to represent the tangent plane $T_{x}\mathcal{T}$ at a point $x\in\mathcal{T}$ as the kernel of a non-zero covector $\alpha$,
$T_{x}\mathcal{T}=\{v\in T_{x}M;\alpha(v)=0\}\,.$
For a torus transverse to $\xi$, such a covector satisfies $\alpha(\xi)\neq 0$; and if $\alpha$ is chosen to be continuous then $\alpha(\xi)$ it is single-signed on the torus. We can multiply $\alpha$ by $-1$ if necessary, to achieve $\alpha(\xi)>0$ everywhere on $\mathcal{T}$. We use such a field of covectors $\alpha$ to represent the tangent planes to $\mathcal{T}$.
Similarly, we can specify a cone at a point as a closed set of non-zero covectors, using the correspondence between the kernel of non-zero covectors and 2D subspaces. As the tangent plane to each invariant torus contains the vector $B$ there, we can restrict attention to cones $C$ consisting of 2D subspaces that contain $B$, equivalently such that $\alpha(B)=0$ for all $\alpha\in C$. Also we can restrict attention to cones $C$ such that $\alpha(\xi)>0$ for all $\alpha\in C$.
With these two restrictions, there is a natural order on covectors. To describe this, it is simplest to introduce another reference vector field. Let $\Omega$ be the volume-form in $M$ and let $\eta$ be a vector field such that $(\eta,\xi,B)$ form a positively oriented frame, i.e. $\Omega(\eta,\xi,B)>0$. For example, one can take $\eta=\xi\times B$. We call $\eta$ a horizontal field. The purpose of $\eta$ is to supplement $\xi$ and $B$, providing a reference frame to define the “slope” of the tangent plane to a torus or of a covector satisfying the above two conditions. Later, we will examine the conefield in a transverse section $\Sigma$ to $B$ (e.g. a poloidal section), so it is convenient to choose $\xi$ and $\eta$ to be tangent to $\Sigma$, which can be achieved by adding suitable multiples of $B$ to each.
For a covector $\alpha$ with $\alpha(B)=0$ and $\alpha(\xi)>0$, we define its slope by
@@
For covectors $\alpha^{\pm}$ with $\alpha^{\pm}(B)=0$, $\alpha^{\pm}(\xi)>0$, we say $\alpha^{+}\geq\alpha^{-}$ if $\sigma^{+}\geq\sigma^{-}$. This condition is independent of multiplying $\alpha$ by a non-zero factor, and of the choice of $\eta$ (as long as the same $\eta$ is used to quantify the slopes of the two covectors and it satisfies the positively oriented frame condition). We will consider cones of the form $\{\alpha:\alpha^{-}\leq\alpha\leq\alpha^{+}\}$. See Figure 1.
Figure 1: A cone, consisting of all the planes like the blue one that lie in the sector between the $\alpha^{+}$ and $\alpha^{-}$ planes not containing $\xi$.
Now, we will construct $\alpha^{\pm}$ iteratively such that if there is an invariant torus $\mathcal{T}$ with tangent plane given by $\alpha$ at the given point then at all stages $\alpha^{-}<\alpha$ and $\alpha<\alpha^{+}$. That is, $\alpha^{\pm}$ provide upper and lower bounds on the slope of any invariant torus of the given class through the given point. If at any stage our construction gives $\alpha^{-}\geq\alpha^{+}$, then we deduce that no such $\alpha$ exists, and hence, there is no invariant torus of the given class through the given point. This will give our primary non-existence condition.
@@
Let $\alpha_{0}$ be a 1-form defined by
$\alpha_{0}=i_{B}i_{\xi}\Omega\,,$
equivalently, $\alpha_{0}=(\xi\times B)^{\flat}$, and note that $\xi$ transverse to $\mathcal{T}$ implies $\alpha_{0}\neq 0$ at all points of $\mathcal{T}$. Then compute ${\alpha}_{t}=\phi_{t}^{*}(\alpha_{0})$ for $t\in[-\tau,\tau]$, some $\tau>0$. This can be done by integrating the adjoint linearised equation about the field line, but it is equivalent to compute the matrix solution of the linearised fieldline flow and apply that to the covector.
Note that $\alpha_{t}(B)=0$ for all $t$. Also, $\alpha_{0}(\xi)=0$. If $\alpha_{t}(\xi)$ happens to be zero for all $t$ then we will deduce nothing, but that is a very special (“shearless”) situation. Let $T_{\pm}$ be the sets of times $t$ for which $\alpha_{t}(\xi)$ is positive and negative, respectively. Let $\sigma^{+}$ be the minimum of the slopes of $\alpha_{t}$ for $t\in T_{+}$ and, $\sigma^{-}$ be the maximum of the slopes of $-\alpha_{t}$ for $t\in T_{-}$. If there is an invariant torus of a given class through the given point, then its slope $\sigma$ satisfies
@@
If the case $\sigma^{+}\leq\sigma^{-}$ occurs during the computation, it will imply that such an invariant torus does not exist. This is a reformulation of the Converse KAM method described in [5]. The advantage of the reformulation is that for points that are not eliminated we still get some information, namely bounds on the slopes of possible invariant tori through them.
The sets $T^{\pm}$ depend on the “shear” of the field $B$ relative to $\xi$. We say $B$ has positive shear relative to $\xi$ if $\frac{d}{dt}(\alpha_{t}\xi)>0$ and negative shear if $\frac{d}{dt}(\alpha_{t}\xi)<0$. In cases of positive shear, then $\alpha_{t}(\xi)>0$ for all small negative $t$, and negative for all small positive $t$. Furthermore, if the shear remains positive along the fieldline then the slope of $\alpha_{t}$ decreases from +$\infty$+ ∞ as $t$ decreases from 00, as long as $\alpha_{t}(\xi)$ remains positive. Similarly, the slope of $-\alpha_{t}$ increases from $\infty$- ∞ as $t$ increases from 00, as long as $-\alpha_{t}(\xi)$ remains positive. If either of ±$\pm\alpha_{t_{\pm}}(\xi)$ reaches 00 as $t_{+}$ decreases or $t_{-}$ increases, then the slopes will already have failed to satisfy $\sigma^{-}<\sigma^{+}$ and one can stop integrating. So the maximum or minimum slopes are achieved for the longest times computed, simplifying the computation. The same applies with obvious changes of sign for negative shear. An advantage of the general procedure above, however, is that it can be applied to cases without definite shear. The shear might be positive in some regions, negative in others, and the fieldline might move from one to the other.
@@
For a magnetic field $B$, direction field $\xi$ and horizontal field $\eta$, given a point $x_{0}=x(0)$ and time $t>0$, pullback the covectors $\alpha_{0}=i_{\xi}i_{B}\Omega$ from $x(\pm t)$ to $x_{0}$ along the fieldline flow and compute the resulting slopes $\sigma^{\pm}$. If there is a time $t$ such that $\sigma^{+}<\sigma^{-}$, then there is no invariant torus for $B$ through $x_{0}$ transverse to $\xi$.
@@
    Compute $\alpha_{t}(\xi)$ and $\alpha_{t}(\eta)$ (note that for any vector $w$ at $x_{0}$, $\alpha_{t}(w)=\alpha_{0}\,M(t)\,w=\Omega(\xi,B,M(t)w)$, where $\xi$ and $B$ are computed at $x(t)$).
@@
    Find the sets $T^{\pm}$ where $\alpha_{t}(\xi)$ is positive/negative at the chosen $x_{0}$.
@@
    Find the the minimum, $\sigma^{+}$, of $\alpha_{t}(\eta)/\alpha_{t}(\xi)$ over $t\in T^{+}$, and the maximum, $\sigma^{-}$, of $\alpha_{t}(\eta)/\alpha_{t}(\xi)$ over $t\in T^{+}$.
@@
For clarity in the exposition and its use in the next subsection, it is convenient to transform the slope $\sigma$, defined relative to the direction field $\xi$ and horizontal field $\eta$, into a slope $m$ in a reference coordinate system (e.g., $(R,z)$ for a poloidal plane $\phi=\phi_{0}$). Assuming that $\xi$ and $\eta$ are tangent to the chosen poloidal plane, we can express the projections of $\alpha$ in components as $\alpha(\xi)=\alpha_{R}\xi^{R}+\alpha_{z}\xi^{z}$ and $\alpha(\eta)=\alpha_{R}\eta^{R}+\alpha_{z}\eta^{z}$, and solve for $\{\alpha_{R},\alpha_{z}\}$, as $\xi$ and $\eta$ are linearly independent by construction. Recalling that $(\alpha_{R},\alpha_{z})$ are the covariant components of a covector, the transformed slope is simply given by $m=-{\alpha_{R}}/{\alpha_{z}}$, that is,

### v2

@@ Abstract
The conefield variant of a Converse KAM method for 3D vector fields, identifying regions through which no invariant 2-tori pass transverse to a specified direction field, is tested on some helical perturbations of an axisymmetric magnetic field in toroidal geometry. This implementation computes bounds on the slopes of invariant tori of a given class and allows to apply a subsidiary criterion to extend the non-existence region, saving significant computation time. The method finds regions corresponding to magnetic islands and chaos for the fieldline flow.
Keywords: Magnetic field, Flux surface, Converse KAM method, Conefield, Killends
@@ 1 Introduction
The present work presents a new application of Converse KAM theory to magnetic fields, complementing the results of [5]. First, it applies an alternative, yet equivalent, implementation of the method described in [6], called the “conefield method”. This version of the Converse KAM method constructs bounds at any point for the slope of the tangent plane to any possible invariant torus through it that is transverse to a prescribed direction field. Secondly, as discussed in the following sections, this approach enables processing the method’s output data to further extend the region of non-existence by a subsidiary criterion called “killends”, inspired by [3].
The paper illustrates the application of the method on the same magnetic fields in toroidal configurations considered in [5], with the same purpose of identifying regions where no invariant tori (flux surfaces) of a given class are present. A class of tori is defined by prescribing a direction field almost everywhere and requiring that the tori be transverse to this field. The primary choice for the direction field is the gradient of an appropriate squared distance function from a closed field line (magnetic axis), with respect to a given metric.
In Section 2, we introduce the magnetic fields tested in this paper. Section 3 outlines the application of the Converse KAM method to magnetic fields and presents the subsidiary criterion to enlarge the detected non-existence region. If preferred, Sections 2 and 3 may be read in the opposite order. In Section 4, we present the results of the numerical implementation of the Converse KAM method and the subsidiary criterion on the chosen magnetic fields. Section 5 presents a discussion of the results and section 6 summarises our conclusions.
@@ 2 Toroidal helical magnetic fields
Given a coordinate system $(x^{1},x^{2},x^{3})$, we will write a magnetic field $B$ in terms of its contravariant components $B^{i}$, rather than its physical ones. This choice reduces the equations of motion for fieldline flow to simply $\dot{x}^{i}(t)=B^{i}(x(t))$; here “time” is to be understood to represent flow along the magnetic field lines at rate $|B|$.
The fields considered have a circular magnetic axis of radius $R_{0}>0$ in the horizontal plane $z=0$, which makes it easy to identify the “principal” class of tori that surround it. They are best described and treated in an adapted toroidal coordinate system $(\psi,\vartheta,\phi)$, which is a variant of the standard toroidal coordinates $(r,\theta,\phi)$. The latter coordinates are related to Cartesian coordinates $(x,y,z)$ through
@@
Following [7], the adapted coordinates $(\psi,\vartheta)$ make the restriction $\beta_{T}$ of the magnetic flux-form $\beta$ to a poloidal section (ϕ=italic-$\phi=$ constant) take the form
@@
where $\wedge$ denotes the exterior product of differential forms (see [8] for a tutorial), where $\psi$ is the toroidal magnetic flux across the poloidal disk of radius $r$ about a point of the magnetic axis, divided by $2\pi$. Thus,
$\beta_{\text{T}}=rRB^{\phi}\,dr\wedge d\theta.$ (2)
For simplicity, we choose our magnetic fields to have $B^{\phi}={B_{0}R_{0}}/{R^{2}}$, which corresponds to a vertical current along the $z$-axis. Then $\vartheta$ can be constructed by equating (1) and (2), up to choice of origin that we take to be at $\theta=0$. Given the previous choices, we arrive at [7, 9, 5]
@@ 2.1 Magnetic fields studied
where $w_{1}\in\mathbb{R}$, $w_{2}\neq 0$, $m,n$ are integers with $m\geq 2$, $f_{mn}$ are smooth functions and $\zeta_{mn}$ arbitrary phases. The factor $\psi^{m/2}$ is to make the resulting vector potential smooth at the coordinate singularity $\psi=0$.
@@ 3 Converse KAM for magnetic fields
Let $B$ be a nowhere-zero $C^{1}$ vector field on an orientable $3D$ manifold $M$. In this paper, a torus in $M$ means a diffeomorphic image of the standard 2-torus $\mathbb{T}^{2}$, with $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ (relaxation to Lipschitz tori will be discussed in section 5). We define a class of tori in $M$ by specifying a $C^{1}$ vector field $\xi$, called “direction field”, and requiring the tori to be transverse to $\xi$; this means that $\xi$ is nowhere tangent to the torus.
In [5], the Converse KAM criterion consisted in pushing forward the direction field $\xi_{x}$ from a point $x$ along $B$ (in one direction of time) and testing whether it becomes a negative multiple of $\xi$ modulo $B$ at some time (i.e. $\exists\,t,\mu\in\mathbb{R},\lambda<0,$ such that $\phi_{t*}\xi_{x}=(\lambda\xi+\mu B)_{\phi_{t}x}$, where $\phi$ is the fieldline flow and $\phi_{t*}$ is its derivative at time $t$). In the present work, we implement a different yet equivalent approach, which we term the ‘conefield formulation’. In this formulation, we pull back a 1-form (in both directions of time) to produce upper and lower bounds on the “slope” of invariant tori through any point; thus, for each point tested, we construct a cone of possible slopes, forming a “conefield”. If at some stage the upper bound is below the lower bound then the set of possible slopes is empty, so there is no invariant torus through that point.
@@ 3.2 Conefield construction and nonexistence condition
We describe here the idea and construction of conefields for the class of invariant tori of a 3D volume-preserving vector field $B$ that are transverse to a direction field $\xi$, by adapting the construction for area-preserving twist maps in [3]. See also the brief description in [6].
If there is a (differentiable) invariant 2-torus $\mathcal{T}$ for $B$ then one can represent the tangent plane $T_{x}\mathcal{T}$ at a point $x\in\mathcal{T}$ as the kernel of a non-zero covector $\alpha_{x}$,
$T_{x}\mathcal{T}=\ker\alpha_{x}=\{v\in T_{x}M;\alpha_{x}(v)=0\}\,.$
In ordinary speak, $\alpha_{x}(v)=n_{x}\cdot v$ where $n_{x}$ is a normal vector to the tangent plane at $x$; but there is no need to invoke a Riemannian metric so we do not express it that way. A covector is a linear functional on a vector space and the set of covectors at a point of a manifold $M$ is called the cotangent space $T^{*}M$.
Multiplication of $\alpha_{x}$ by any non-zero scalar produces the same kernel, so one could consider $\alpha_{x}$ to be in the quotient by non-zero scalars of the cotangent space minus the origin. It is convenient, however, to assign a “co-orientation” to a 2D subspace in 3D, i.e. to say one side of it is positive and the other negative. This can be done by the sign of $\alpha_{x}$ off its kernel. So we consider $\alpha_{x}$ to be in the quotient of $T_{x}^{*}M\setminus\{0\}$ by only positive scalars.
Consistently using the subscript x to indicate evaluation at a point $x$ will lead to excessive notation when we later introduce time-dependent covectors and add superscripts $\pm$. Therefore, we will omit the subscript when the context makes it clear.
For a torus transverse to $\xi$, such a covector satisfies $\alpha(\xi)\neq 0$. A natural co-orientation is induced by taking $\alpha(\xi)>0$, corresponding to saying $\xi$ is on the positive side of $\ker\alpha$.
The tangent plane to an invariant torus at a point $x$ contains the vector $B_{x}$, so we can restrict attention to covectors $\alpha$ such that $\alpha(B)=0$.
For a covector $\alpha$ satisfying both $\alpha(\xi)>0$ and $\alpha(B)=0$, we can define its “slope”, corresponding to a notion of slope of the plane defined by its kernel. To do this, we introduce another reference vector field $\eta$. Let $\Omega$ be the volume-form in $M$ (that is preserved by $B$) and let $\eta$ be a vector field such that $(\eta,\xi,B)$ form a positively oriented frame, i.e. $\Omega(\eta,\xi,B)>0$. For example, one can take $\eta=\xi\times B$ (though when we choose a transverse section to $B$ later, it will be convenient to choose both $\xi$ and $\eta$ to be tangent to the section by subtracting off suitable multiples of $B$). We call $\eta$ a horizontal field. Then we define the slope $\sigma\in\mathbb{R}$ of such a covector $\alpha$ by
@@
The justification for the name is that $\alpha(\eta+\sigma\xi)=0$ so $\eta+\sigma\xi\in\ker\alpha$; this vector has slope $\sigma$ in coordinates with horizontal unit vector $-\eta$ and vertical unit vector $\xi$ (perhaps it would be considered more natural to take the opposite sign for $\eta$ and define $\sigma$ with a minus sign). We say one such covector $\alpha^{+}$ is greater than or equal to another $\alpha^{-}$ iff its slope is greater than or equal to the slope of the other; we write this as $\alpha^{+}\succeq\alpha^{-}$. The same definition can be applied to covectors with $\alpha(\xi)<0,\alpha(B)=0$, but it is important to realise that if $\alpha^{+}\succeq\alpha^{-}$ then also $-\alpha^{+}\succeq-\alpha^{-}$.
We extend the definition of the order on covectors $\alpha$ with $\alpha(B)=0$ to the multiples of
$\alpha_{0}=i_{B}i_{\xi}\Omega$ (7)
[…]
The conefield method constructs bounds on the slopes of all invariant tori that are transverse to $\xi$. A cone at a point $x\in M$ is a set $C_{x}$ of non-zero covectors at $x$; in practice, we will take cones of the form $\alpha_{x}^{-}\preceq\alpha_{x}\preceq\alpha_{x}^{+}$ for some $\alpha_{x}^{\pm}$ (see Figure 1).
Figure 1: A cone, consisting of all the covectors at a point whose kernels are planes, like the blue one, that lie in the sector between those for $\alpha^{+}$ and $\alpha^{-}$ that does not contain $\xi$.
A conefield $C$ is a choice of cone at each point of $M$ (to allow the killends extension, we will require $C_{x}$ to be measurable with respect to $x$). We say a conefield bounds the slopes of all invariant tori transverse to $\xi$ if for all points $x$ on such an invariant torus the cone $C_{x}$ contains a covector representing the tangent plane to the torus there.
We will start with the conefield $C$ defined by $\alpha^{+}=\alpha_{0},\alpha^{-}=-\alpha_{0}$. It trivially bounds the slopes of all invariant tori of the given class, because it contains all possible slopes at all points. Indeed, by transversality, we have the strict order $-\alpha_{0}\prec\alpha\prec\alpha_{0}$ for any covector $\alpha$ representing the tangent plane to an invariant torus of the given class. But after an initial stage of an iteration that we will describe, we will typically obtain that $\alpha^{\pm}(\xi)>0$ at all points.
Now we will modify $\alpha^{\pm}$ iteratively such that if there is an invariant torus $\mathcal{T}$ with tangent plane given by $\ker\alpha$ at the given point then at all stages $\alpha^{-}\prec\alpha$ and $\alpha\prec\alpha^{+}$. That is, $\alpha^{\pm}$ provide (strict) upper and lower bounds on the slope of any invariant torus of the given class through the given point. If at any stage our construction gives $\alpha^{-}\succeq\alpha^{+}$, then we deduce that no such $\alpha$ exists, and hence, there is no invariant torus of the given class through the given point. This will give our primary non-existence condition.
@@
For given choice of direction field $\xi$, let $\alpha_{0}$ be the 1-form defined in (7), equivalently, $\alpha_{0}=(\xi\times B)^{\flat}$, and note that $\xi$ transverse to $\mathcal{T}$ implies $\alpha_{0}\neq 0$ at all points of $\mathcal{T}$ ($\alpha_{0}\neq 0$ means there exists a vector $v$ on which $\alpha_{0}(v)\neq 0$, e.g. $v=\xi\times B$). Then compute the pullback
${\alpha}_{t}=\phi_{t}^{*}(\alpha_{0})$
for $t\in[-\tau,\tau]$, some $\tau>0$. This can be done by integrating the adjoint linearised equation about the field line, but it is equivalent to compute the matrix solution of the linearised fieldline flow and apply that to the covector (as we will spell out below).
Note that $\alpha_{t}(B)=0$ for all $t$. Also, $\alpha_{0}(\xi)=0$. If $\alpha_{t}(\xi)$ happens to be zero for all $t$ then we will deduce nothing, but that is a very special (“shearless”) situation. Let $T_{\pm}$ be the sets of times $t$ for which $\alpha_{t}(\xi)$ is positive and negative, respectively. Let $\sigma^{+}$ be the minimum of the slopes of $\alpha_{t}$ for $t\in T_{+}$, and $\sigma^{-}$ be the maximum of the slopes of $-\alpha_{t}$ for $t\in T_{-}$ (the slope of $-\alpha_{t}$ is the same as for $\alpha_{t}$ but we include the minus sign because the resulting $\alpha^{-}$ that we want to use is $-\alpha_{t}$ for the maximising $t\in T_{-}$). If there is an invariant torus of a given class through the given point, then its slope $\sigma$ satisfies
@@
We define the resulting cone by $\alpha^{+}=\alpha_{t}$ for the minimising value of $t\in T_{+}$ and $\alpha^{-}=-\alpha_{t}$ for the maximising value of $t\in T_{-}$.
As $\tau$ is increased, the cone at a point can only get tighter. If the case $\sigma^{+}\leq\sigma^{-}$ occurs during the computation, it will imply that such an invariant torus $\mathcal{T}$ does not exist. This is a reformulation of the Converse KAM method described in [5]. The advantage of the reformulation is that for points that are not eliminated we still get some information, namely bounds on the slopes of possible invariant tori through them.
The sets $T_{\pm}$ depend on the “shear” of the field $B$ relative to $\xi$. We say $B$ has positive shear relative to $\xi$ if $\frac{d}{dt}(\alpha_{t}\xi)<0$ at $t=0$ and negative shear if $\frac{d}{dt}(\alpha_{t}\xi)>0$ (this looks backwards, but $\alpha_{t}$ is the pullback of $\alpha_{0}$). In cases of positive shear, then $\alpha_{t}(\xi)>0$ for all small negative $t$, and negative for all small positive $t$. Furthermore, if the shear remains positive along the fieldline then the slope of $\alpha_{t}$ decreases from +$\infty$+ ∞ as $t$ decreases from 00, as long as $\alpha_{t}(\xi)$ remains positive. Similarly, the slope of $-\alpha_{t}$ increases from $\infty$- ∞ as $t$ increases from 00, as long as $-\alpha_{t}(\xi)$ remains positive. So in this case the set $T_{+}$ consists of negative times and the set $T_{-}$ consists of positive times. Denoting $t_{+}$ the backwards integration time and $t_{-}$ the forwards integration time, if either of ±$\pm\alpha_{t_{\pm}}(\xi)$ reaches 00 as $t_{+}$ decreases or $t_{-}$ increases, then the slopes will already have failed to satisfy $\sigma^{-}<\sigma^{+}$ and one can stop integrating. So, in the case of positive shear the maximum or minimum slopes are achieved for the longest (backwards and forwards, respectively) times computed, simplifying the computation. The same applies with obvious changes of sign for negative shear. An advantage of the general procedure above, however, is that it can be applied to cases without definite shear. The shear might be positive in some regions, negative in others, and the fieldline might move from one to the other.
@@
For a magnetic field $B$, direction field $\xi$ and horizontal field $\eta$, given a point $x_{0}=x(0)$ and time $\tau>0$, let $\alpha_{t}$ be the pullback of the covector $\alpha_{0}=i_{B}i_{\xi}\Omega$ from $x(t)$ to $x_{0}$ along the fieldline flow for all $t\in[-\tau,\tau]$ and compute the resulting slope $\sigma_{t}$. If there are times $t_{\pm}$ such that $\sigma_{t^{+}}\leq\sigma_{t^{-}}$ and ±$\pm\alpha_{t_{\pm}}(\xi)>0$, then there is no invariant torus for $B$ through $x_{0}$ transverse to $\xi$.
In this statement, we didn’t insist on finding the maximum and minimum of $\sigma_{t}$ over $t\in T_{\pm}$, because the extra freedom is useful in practice.
@@
    Compute $\alpha_{t}(\xi)$ and $\alpha_{t}(\eta)$, using that for any vector $w$ at $x_{0}$, $\alpha_{t}(w)=\alpha_{0}\,M(t)\,w=\Omega(\xi,B,M(t)w)$, where $\Omega$, $\xi$ and $B$ are computed at $x(t)$.
@@
    Find the sets $T_{\pm}$ of $t$ where $\alpha_{t}(\xi)$ is computed to be positive/negative at the chosen $x_{0}$.
@@
    Find the minimum, $\sigma^{+}$, of $\alpha_{t}(\eta)/\alpha_{t}(\xi)$ over $t\in T_{+}$, and the maximum, $\sigma^{-}$, of $\alpha_{t}(\eta)/\alpha_{t}(\xi)$ over $t\in T_{-}$.
@@
  6. 6.
    Define $\alpha_{\tau}^{+}$ to be $\alpha_{t}$ at the maximising time $t\in T_{+}$ and $\alpha_{\tau}^{-}$ to be $-\alpha_{t}$ at the minimising time $t\in T_{-}$, and $\sigma^{\pm}_{\tau}$ the corresponding slopes.
Figure 2 illustrates the typical evolution of $\alpha_{\tau}^{\pm}(\eta)$ and $\alpha_{\tau}^{\pm}(\xi)$, along with the associated slopes $\sigma_{\tau}^{\pm}$, as functions of time $\tau$ for two initial conditions for the example magnetic fields that we study here, [^2] The specific parameter values of the two simulations correspond to (13), (14) with $\varepsilon_{21}=\varepsilon_{32}=0.003$. which have positive shear $(w_{2}>0)$. The left column, corresponding to a point outside the non-existence region, shows how the cone at the point becomes narrower as time evolves, with the slopes approaching a common value. The right column, corresponding to a point in the non-existence region, shows how the computation is stopped when $\sigma_{\tau}^{+}<\sigma_{\tau}^{-}$ (to make the crossing clear, we integrated a bit further than necessary). Note that the integration time is much shorter for the second case. Note also that on an orbit with positive Lyapunov exponent, $\alpha_{t}$ can be expected to grow at the Lyapunov rate, but the power of the Converse KAM method is that typically, the greater the Lyapunov exponent, the shorter the time it takes to reach a decision.
Figure 2: Typical evolution of the values of $\alpha_{\tau}^{\pm}(\eta)$ and $\alpha_{\tau}^{\pm}(\xi)$ (top), along with the associated slopes $\sigma_{\tau}^{\pm}$ (bottom) for the magnetic field derived from (14), for two different initial conditions: [(a)(b)] a point in the complement of the non-existence region ($(\tilde{y}_{1},\tilde{z}_{1})=(0.3,0.2)$); and [(c)(d)] a point inside the non-existence region ($(\tilde{y}_{2},\tilde{z}_{2})=(0.35,0.45)$). The insets show blowups of the indicated regions.
For use in the next subsection, it is convenient to transform the slope $\sigma$, defined relative to the direction field $\xi$ and horizontal field $\eta$, into a slope $m$ in a reference coordinate system (e.g., $(R,z)$ for a poloidal plane $\phi=\phi_{0}$). Assuming that $\xi$ and $\eta$ are tangent to the chosen poloidal plane, we can express the projections of $\alpha$ in components as $\alpha(\xi)=\alpha_{R}\xi^{R}+\alpha_{z}\xi^{z}$ and $\alpha(\eta)=\alpha_{R}\eta^{R}+\alpha_{z}\eta^{z}$, and solve for $\{\alpha_{R},\alpha_{z}\}$, as $\xi$ and $\eta$ are linearly independent by construction. Recalling that $(\alpha_{R},\alpha_{z})$ are the covariant components of a covector, the transformed slope is simply given by $m=-{\alpha_{R}}/{\alpha_{z}}$, that is,